Please help with multi step question- 1 question with 6 parts!!
The table below gives the number of hours seven randomly selected students spent studying and their corresponding midterm exam grades. Using this data, consider the equation of the regression line, y bo + bix, for predicting the midterm exam grade that a student will earn based on the number of hours spent studying. Keep in mind, the correlation coefficient may or may not be statistically significant for the data given. Remember, in practice, it would not be appropriate to use the regression line to make a prediction if the correlation coefficient is not statistically significant. Hours Studying 0.5 2.5 3 4 4.5 5 5.5 Midterm Grades62 74 86 87 97 98 100 Table Copy Data Step 1 of 6: Find the estimated slope. Round your answer to three decimal places. AnswerHow to Enter) 2 Points KeypadTables Nex Prev
he table below gives the number of hours seven randomly selected students spent studying and their corresponding nidterm exam grades. Using this data, consider the equation of the regression line, y- bo + bjx, for predicting the midterm exam grade that a student will earn based on the number of hours spent studying Keep in mind, the correlation coefficient may or may not be statistically significant for the data given. Remember, in practice, it would not be appropriate to use the regression line to make a prediction if the correlation coefficient is not statistically significant. Hours Studying 0.5 2.5 3 4 4.5 55.5 Midterm Grades 62 74 86 87 97 98 100 Table Copy Data Step 2 of 6: Find the estimated y-intercept. Round your answer to three decimal places Answer(How to Enter) 2 Points KeypadTables
The table below gives the number of hours seven randomly selected students spent studying and their corresponding midterm exam grades. Using this data, consider the equation of the regression line, bobix, for predicting the midterm exam grade that a student will earn based on the number of hours spent studying. Keep in mind, the correlation coefficient may or may not be statistically significant for the data given. Remember, in practice, it would not be appropriate to use the regression line to make a prediction if the correlation coefficient is not statistically significant. Hours Studying 0.5 2.5 3 4 4.5 5 5.5 Midterm Grades 62 74 86 87 97 98 100 Table Copy Data Step 3 of 6: Find the estimated value of y when x - 4.5. Round your answer to three decimal places. Answer(How to Enter) 2 Points Keypad Tables
The table below gives the number of hours seven randomly selected students spent studying and their corresponding midterm exam grades. Using this data, consider the equation of the regression line) bo + bix, for predicting the midterm exam grade that a student will earn based on the number of hours spent studying, Keep in mind, the correlation coefficient may or may not be statistically significant for the data given. Remember, in practice, it would not be appropriate to use the regression line to make a prediction if the correlation coefficient is not statistically significant. Hours Studying 0.5 2.5 3 4 4.5 5 5.5 Midterm Grades 62 74 86 8797 98 100 Table Copy Data Step 4 of 6: According to the estimated linear model, if the value of the independent variable is increased by one unit, then the change in the dependent variable ý is given by? Answer 2 Points Keypad Tables ev O bo Next Oy
Step 4 of 6: According to the estimated linear model, if the value of the independent variable is increased by one unit, then the change in the dependent variable ý is given by? Answer 2 Points Keypad Tables O bo ox
media%2F0c8%2F0c882456-a0cd-4954-b392-dc
Tabl Copy Data Step 6 of &: Find the value of the coefficient of determination. Round your answer to three decimal places Answer(How to Enter) 2 Points d Keypad Tables
Answer #1

Solution:

We have formulas given as below:

Slope = b1 = (∑XY – n*Xbar*Ybar)/(∑X^2 – n*Xbar^2)

y-intercept = b0 = Ybar – b*Xbar

Calculation table is given as below:

No.

X

Y

X^2

Y^2

XY

1

0.5

62

0.25

3844

31

2

2.5

74

6.25

5476

185

3

3

86

9

7396

258

4

4

87

16

7569

348

5

4.5

97

20.25

9409

436.5

6

5

98

25

9604

490

7

5.5

100

30.25

10000

550

Total

25

604

107

53298

2298.5

Xbar = ∑X/n = 25/7 = 3.571429

Ybar = ∑Y/n = 604/7 = 86.28571

b1 = (∑XY – n*Xbar*Ybar)/(∑X^2 – n*Xbar^2)

b1 = (2298.5 - 7*3.571429*86.28571)/( 107- 7*3.571429^2)

b1 = 7.97984

Slope = 7.97984

b0 = Ybar – b*Xbar

b0 = 86.28571 - 7.97984*3.571429

b0 = 57.78628

y-intercept = 57.78628

Correlation coefficient = r = [n∑xy - ∑x∑y]/sqrt[(n∑x^2 – (∑x)^2)*(n∑y^2 – (∑y)^2)]

r = [7*2298.5- 25*604]/sqrt[(7*107– 25^2)*(7*53298– 604^2)]

r = 0.97713

Coefficient of determination = r^2 = 0.97713*0.97713 = 0.954783

Answers:

Step 1

Answer:

Slope = 7.980

Step 2

Answer:

y-intercept = 57.786

Step 3

Answer: 93.696

We are given X = 4.5

We have

Y = 57.786 + 7.980*X

Y = 57.786 + 7.980*4.5

Y = 93.696

Step 4

Answer: b1

Step 5

Answer: 3.304

Error = Residual = Observed – predicted

When x = 4.5

Observed = 97

Predicted = Y = 57.786 + 7.980*X

Predicted =Y = 57.786 + 7.980*4.5

Predicted = 93.696

Error = Residual = 97 - 93.696

Error = Residual = 3.304

Step 6

Answer: 0.955